A basketball player makes baskets from the foul line 88% of the time. If she makes 8 attempts, what is the probability she will miss no more than twice?
Answer #1

since this is the problem of binomial distribution where we have given

p=0.88 and q =1-p =1-0.88=0.12

n=8 and r=2

general binomial formula is P(x=r) =nCrpr(1-p)n-r

=P(x<2) = 8C0(.88)0(1-0.88)8 + 8C1(.88)1(1-0.88)7 =2.56 *10-6